OpenStudy (anonymous):

I got help for this problem before, but just checked and it's not correct. What am I doing wrong with this problem? f(x)=3sqrt(5-x)+3 y=3sqrt(5-x)+3 3sqrt(5-x)=y-3 (3sqrt(5-x)^3=(y-3)^3 5-x=(y-3)^3 x=-(y-3)^3-5

OpenStudy (anonymous):

y = 3rt(5-x) + 3 ; i assume you wanna invert it x = (5-y)^(1/3) + 3 x-3 = (5-y)^(1/3) (x-3)^3 = 5-y (x-3)^3 -5 = -y -(x-3)^3 +5 = y

OpenStudy (anonymous):

last step

OpenStudy (anonymous):

This is the help I got, but it's still not correct.

OpenStudy (anonymous):

y = 3rt(5-x) + 3 ; x = (5-y)^(1/3) + 3 These two are not same! even not inverses of each other!

OpenStudy (anonymous):

I was confused with that. that's what the other tutor person on here sent me.

OpenStudy (amistre64):

tutor? :)

OpenStudy (anonymous):

question answerer?

OpenStudy (amistre64):

free helper that is prone to error ;)

OpenStudy (anonymous):

I tried to send you a message but you werent on anymore

OpenStudy (anonymous):

You're fine to be wrong, I just want to know what I'm doing.

OpenStudy (anonymous):

Inverse of f(x)=3rt(5-x)+3 is g(x)=5-(x-3)^3

OpenStudy (anonymous):

sorry if you felt I was blaiming

OpenStudy (amistre64):

no feelings hurt :)

OpenStudy (anonymous):

I put that and it told me that there was a syntax error

OpenStudy (amistre64):

y = cbrt(5-x) + 3 ; we can usually find the inverse by swapping x and y ; and resolving for y. x = cbrt(5-y) +3 x-3 = cbrt(5-y) ; ^3 each side... (x-3)^3 = 5-y ; +y and -(x-3)^3 to both sides y = 5 -(x-3)^3 which is exaclty the same as: y = -(x-3)^3 + 5 which sauhkik says aint the solution lol

OpenStudy (amistre64):

cbrt(5-y) = (5-y)^(1/3) they are the exact same thing..

OpenStudy (anonymous):

Ok, I tried it again and it was happy with it. I must have typed that one wrong or something.

OpenStudy (anonymous):

Yeah I know that's the same thing.

OpenStudy (amistre64):

:) im glad it worked for you :)

OpenStudy (anonymous):

:) this online hmwk may be the end of me..... :p

OpenStudy (amistre64):

programs can be sticklers for "do it exactly like i say" mentality lol

OpenStudy (anonymous):

It also gets unhappy when simplifying raticals... took me forever figure that one out.

OpenStudy (amistre64):

have you tried yelling at it? :)

OpenStudy (anonymous):

haha, yeah that has happend. the little thinking dudes in my computer dont seem to hear me....